OpenStudy (anonymous):

A certain liquid has a vapor pressure of 92.0 Torr at 23.0 °C and 313.0 Torr at 45.0 °C. Calculate the value of ΔH°vap for this liquid

OpenStudy (anonymous):

Calculate the normal boiling point of this liquid.

OpenStudy (anonymous):

Do you know the formula?

OpenStudy (anonymous):

yea, its the Clausius-Clapeyron equation

OpenStudy (anonymous):

i know how to get the first part, its the temperature i dont know how to get

OpenStudy (anonymous):

Well show me what you know and we'll go from there.

OpenStudy (anonymous):

ln(p2)-ln(p1)+(Hvap/R)(1/T1-1/T2)

OpenStudy (anonymous):

rearrange for Hvap

OpenStudy (anonymous):

ln(p2/p1)-(1

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

ln(p2/p1)-(1/T1-1/T2)/R=Hvap

OpenStudy (anonymous):

I think thats wrong

OpenStudy (anonymous):

I get 43.55 as H vap

OpenStudy (anonymous):

i re- arranged it wrong

OpenStudy (anonymous):

Here: logP2/P1= delta H/(2.303*R) * (1/T1-1/T2) P1=92 torr P2=372, T1=296k, T2=318 R=1.99 cal/mole-k log372/92=0.6067 2.303*R=2.303*1.99=4.5829cal/mole-k ... 1/T1-1/T2=0.0002337K deltaH = 0.6067x4.5829/0.0002337 =11897 cal 11897/239kj = 49.78kj I believe my math is right.This is the first half of the problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand so far?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Do you know what to do next?

OpenStudy (anonymous):

um nope. thats the part i got stuck on

OpenStudy (anonymous):

log760/92=0.9170 , P1=92 torr 2.303*R=2.303*1.99=4.5829cal/mole-k ... 1/T1-1/T2= (1/296-1/T2) , T1=273+23=296k 1/T1-1/T2=( 0.003378 - 1/T2) Heat of vaporisation = 47.9kj 47.9*239cal = 11448cal 0.9170*4.5829/11448=0.003378-1/T2 0.0003671=0.003378-1/T2 1/T2=0.003378-0.0003871=0.0030109 ... T2=332K or 59 degc I hope you understand!

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

What are your anwser choices, if you have any?

OpenStudy (anonymous):

i dont have any

OpenStudy (anonymous):

ry my H value

OpenStudy (anonymous):

try*

OpenStudy (anonymous):

could u try it with my H value...i try it myself and i get a very big number

OpenStudy (anonymous):

could u help with another question

OpenStudy (anonymous):

Sorry I had my own school work to take care of.